OpenStudy (anonymous):

logb(2)=p, find an expression in terms of p for the following. logb(6)

OpenStudy (whpalmer4):

\[\log_b 2 = p\]means that \[b^p = 2\]

OpenStudy (anonymous):

\[\log _b6=\log_b(2\times 3)=\log _b2+\log _b3=?\]

OpenStudy (anonymous):

p=logb(3)?